# FIPO: Free-form Instruction-oriented Prompt Optimization with Preference Dataset and Modular Fine-tuning Schema

### v1

@@ 4 Experiments
In this section, we begin with our experimental settings, fine-tuning performances, then follow with comparisons with SOTA, and case analysis [^4] Hyperparameters and training cost are in Appendix A..
@@ 4.1 Experimental Settings
Baselines. We compare our results with two SOTA LLM-based discrete APO methods: APE Zhou et al. (2023) and PromptAgent Wang et al. (2023a). APE, stands for Automatic Prompt Engineer, is a template-based strategy that ask one LLM to generate a pool of candidate prompts based on the templates, then select one according to the evaluation scores. PromptAgent eliminates templates and replaces with Monte Carlo Tree Search Abramson (2014) for using a evaluator model to guide the generator. It worth to note that both APE and PromptAgent are training-free, aiming to model-oriented prompt optimization in an ad-hoc manner, with GPT-3.5 and GPT-4 APIs, while we realize completely offline training and inference during the prompt optimization stage. Moreover, since PromptAgent aims to optimized the beginning instruction-agnostic learning sentence, as shown in Figure 1, it requires to attach the following examples and task instruction for few-shot testing. In contrast, APE and our FIPO combine unchanged leading sentence, fixed in-context examples, and optimized task instruction in testing. We use Tulu2 models as our bases, which is a fine-tuned version of Llama2 Touvron et al. (2023) trained on a mix of publicly available datasets Ivison et al. (2023). Three versions of Tulu2 models are released: 7B, 13B and 70B. The “B” here refers to the total number of parameters in unit of billions.
@@
Human 24.77 36.21 73.35 51.17 51.22 47.79
Best 13B Optimizer 25.38($\uparrow$) 35.93 71.23 49.03 51.18 47.36
SFT-70B (L) 26.44($\uparrow$) 32.46 73.90($\uparrow$) 50.64 50.03 46.33
SFT-70B (F) 21.43 32.92 74.39($\uparrow$) 49.97 51.55($\uparrow$) 46.96
DPO-70B (L) 24.24 41.88($\uparrow$) 76.24($\uparrow$) 55.87($\uparrow$) 54.09($\uparrow$) 51.39($\uparrow$)
DPO-70B (F) 27.74(normal-$\uparrow$) 35.56 74.17($\uparrow$) 54.93($\uparrow$) 52.73($\uparrow$) 49.07($\uparrow$)
IPO-70B (L) 22.95 42.50($\uparrow$) 76.95($\uparrow$) 55.97($\uparrow$) 54.17($\uparrow$) 51.58($\uparrow$)
IPO-70B (F) 25.00($\uparrow$) 39.21($\uparrow$) 76.84($\uparrow$) 56.01($\uparrow$) 54.29($\uparrow$) 50.94($\uparrow$)
IPL-DPO-70B (L) 21.96 42.04($\uparrow$) 75.64($\uparrow$) 56.01($\uparrow$) 55.13($\uparrow$) 51.84($\uparrow$)
IPL-DPO-70B (F) 25.13($\uparrow$) 35.25 74.95($\uparrow$) 50.46 52.12($\uparrow$) 48.10($\uparrow$)
IPL-IPO-70B (L) 21.96 42.92(normal-$\uparrow$) 75.37($\uparrow$) 55.64($\uparrow$) 55.03($\uparrow$) 51.94($\uparrow$)
IPL-IPO-70B (F) 26.67($\uparrow$) 39.60($\uparrow$) 77.11(normal-$\uparrow$) 56.71(normal-$\uparrow$) 56.02(normal-$\uparrow$) 52.13(normal-$\uparrow$)
Table 2: Evaluation results of out-of-box Tulu2-7B model. “Human” refers to use raw few-shot prompts written by humans, while others use optimized prompts generated by different FIPO optimizer. The number attached with the benchmark name refers to the number of in-context examples (e.g., BBH (3) stands for 3-shot evaluation on BBH). The “(L)” or “(F)” letter attached with FIPO optimizer refer to parameter-efficient tuning LORA Hu et al. (2022) and full-parameter tuning, respectively. The “X” part of “IPL-X” refers to the different loss style mentioned in Equation 5. More detailed results are report in Table 8 in Appendix D.
Evaluation Benchmarks. We include five popular benchmarks across two most common formats: (1) GSM8k Cobbe et al. (2021), a generative dataset contains 1.3k primary level math questions; (2) BigBenchHard (BBH) Suzgun et al. (2023), which involves 23 challenging reasoning tasks. BBH has 6.4k testing samples, and asks for generative answering; (3) PiQA Bisk et al. (2020), in which 1.8k common physical knowledge questions are proposed, alongside with binary candidate choices; (4) CosmosQA Huang et al. (2019). There are around 3k commonsense-based multi-choice questions in CosmosQA, equipped with four candidate options; (5) MMLU Hendrycks et al. (2021), which is one of the largest multi-choice benchmarks. MMLU covers 14k questions. For our FIPO experiments, we report results on all five benchmarks. While differently, since both APE and PromptAgent only provide evaluations on 6 tasks of BBH, we report the comparison results aligned with their settings. As for the result metrics, as shown in Figure 1, either “generation” or “multi-choice” benchmarks takes few-shot format with strict answering templates (e.g., “The answer is X”). Herein, we are able to report the accuracy score for all benchmarks.
Generation Multi-choice
Llama2-7B Human 8.89 31.21 62.78 43.09 46.58 41.73
Touvron et al. (2023) FIPO Optimizer 11.70 33.50 69.37 52.11 54.56 48.10
Tulu2-13B Human 39.06 36.49 76.62 55.13 57.43 52.53
Ivison et al. (2023) FIPO Optimizer 40.17 40.26 78.58 57.68 59.10 54.79
Baichuan2-13B Human 46.81 37.95 68.56 51.88 57.46 52.36
Yang et al. (2023a) FIPO Optimizer 48.12 39.95 74.77 56.88 58.32 54.35
Table 3: Evaluation results of various downstream generator LLMs.
GPT-3.5 Llama2-7B Tulu2-7B
BBH Tasks Human APE PromptAgent FIPO Human PromptAgent FIPO Human PromptAgent FIPO
Penguins 0.595 0.747 0.797 0.614 0.436 0.426 0.458 0.336 0.470 0.525
Geometry 0.227 0.490 0.670 0.738 0.536 0.544 0.542 0.089 0.350 0.415
Epistemic 0.452 0.708 0.806 0.865 0.611 0.775 0.754 0.581 0.691 0.644
Object Counting 0.612 0.716 0.860 0.660 0.318 0.312 0.350 0.690 0.577 0.637
Temporal 0.720 0.856 0.934 0.756 0.362 0.643 0.683 0.155 0.176 0.210
Causal Judgment 0.470 0.570 0.670 0.759 0.547 0.547 0.616 0.832 0.882 0.891
Average 0.513 0.681 0.790 0.732 0.468 0.541 0.567 0.447 0.524 0.554
Table 4: Comparison between FIPO optimizer and previous methods.
4.2 Experimental Results
Fine-tuning Analysis. We first report main results of testing optimized prompts across downstream benchmarks in Table 2. We fine-tune Tulu2-13B and Tulu2-70B models though different strategies as mentioned above, incorporating with parameter-efficient (e.g., LORA Hu et al. (2022)) and full-parameter way, and then use fine-tuned FIPO optimizer to generate new instruction-oriented prompts for downstream testing. The testing is carried with out-of-box Tulu2-7B model in few-shot format.
@@ 4.2 Experimental Results
We observe several findings: (1) “Small” Tulu2-13B are not up to the difficult prompt optimization task; (2) When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task; (3) When contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%; (4) Optimizer fine-tuned in LORA way achieve better results when combined with DPO, IPO and IPL-DPO strategies. However, full-parameter tuning wins in SFT and IPL-IPO settings; (5) In general, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline. The results indicate the effectiveness of its regularized design in Equation 3; (6) We receive more robust improvements on multi-choice benchmarks. The reason may lie in the output limitations, since evaluations on multi-choice benchmarks are constrained to serial options, while evaluations on generation benchmarks are free to all text.
We further check if our optimized prompts have general gains on different downstream generators in Table 3, and conclude that our optimized instruction-oriented prompts are generally beneficial. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average. The improvements reduced as the generator grows larger. This may be due to the fact that larger LLMs has richer world knowledge, so the enhanced instructions are less necessary.
Analysis on Iterative Preference Learning. To validate the self-rewarding capability of fine-tuning prompt optimizer within our IPL pipeline, we present a further epoch-by-epoch checking in Figure 4. Compared with basic DPO and IPO preference tuning, to pump them into our IPL pipeline can advance to better accuracy expect for the full-parameter tuning of IPL-DPO-70B. It has been visualized that either Lora or full-parameter tunings of DPO under the IPL pipeline brings no self-rewarding gains, showing as marginal fluctuations. However, the IPO fine-tuning fits well. We notice that both Lora and full-parameter tunings of IPO under the IPL pipeline behave raising tendencies, and finally breakthrough the performance ceiling of basic IPO. Particularly, the full-parameter tuning of IPL-IPO-70B exhibits a steeper upward trend.
Figure 4: The tendency visualization of IPL (better view with colors). We denote average accuracy over benchmarks of every checking models as markers. The blue triangles and red triangles refer to IPL-DPO-70B and IPL-IPO-70B models respectively, which is in accordance with the last four lines in Table 2.
Raw Task Instruction:
Calculate the average of the list [12, 34, 56, 75].
Raw Task Response:
Sure, list is a basic data structure in Python. We can compute the average of the list by: (1) list = [12, 34, 56, 75]; (2) average = sum(list) / len(list); (3) print(average). The final answer is sum([12, 34, 56, 75]) / len([12, 34, 56, 75]) = 177 / 4 = 44.25.
Ground Truth Response:
To calculate the average value of a list of numbers, you sum up all the numbers and then divide by the count of numbers in the list. For the list [12, 34, 56, 75], the sum of the numbers is: 12 + 34 + 56 + 75 = 177. There are 4 numbers in the list. Now, divide the sum by the count of numbers to get the average: Average = 177 / 4 = 44.25. So, the average value of the list [12, 34, 56, 75] is 44.25.
Suboptimal Optimized Prompt:
You are an expert in mathematics and someone asks you to calculate the average of a list of numbers. Write a complete solution to explain how to calculate the average step-by-step, using the list [12, 34, 56, 75] as an example. Make sure to provide the final answer in your response.
Optimal Optimized Prompt:
We need to calculate the average of a set of numbers, specifically [12, 34, 56, 75]. Please provide the process in completing the task, including a step-by-step explanation of how to calculate the average, summing all the items in the list, and then dividing this total by the number of elements in the list. Your response should be detailed, clear, and accurate, similar to the manner a math tutor would explain to a student learning this concept for the first time.
Optimized Instruction from SFT-70B (Full)
As an AI language model, your task is to provide a detailed explanation on how to calculate the average of a list of numbers. The list consists of numbers [12, 34, 56, 75]. I would like you to provide a step-by-step guide on how to calculate the average. Please also explain what is meant by “average” in this context, and how the calculation is different from the total sum of the numbers in the list divided by the number of elements in the list. Please start by defining what “average” means in the context of a list of numbers. Explain how this concept relates to the process of calculating the average of a list. Then, provide a step-by-step guide on how to calculate the average of the list [12, 34, 56, 75].
Optimized Instruction from DPO-70B (Full)
Hi AI, could you accurately compute the mean value of this given list of distinct integer numbers? The list consists of: 12, 34, 56, and 75. It would be helpful if you could provide the mathematical formula or steps you took in achieving the solution. Also, make sure to include the final calculated average value in your response.
Optimized Instruction from IPO-70B (Full)
Given the list of integers [12, 34, 56, 75], calculate the average value by finding the sum of all elements and dividing it by the total number of elements in the list. Express the result as a floating-point number up to two decimal places. Show the step-by-step calculation to obtain the average and include a brief explanation of the process.
Optimized Instruction from IPL-IPO-70B (Full)
@@
Table 5: An example from GSM8K benchmark, which is same as the example in Figure 2.
Comparison with SOTA. Except for the effectiveness of fine-tuning strategies, we would like to compare the FIPO research with previous methods. Table 4 reports experimental results on 6 BBH tasks, cross-evaluating with different methods and downstream generators. On the one hand, PromptAgent obtain significant improvements over all tasks, compared with simple few-shot prompts written by humans and optimized prompts generated by APE. However, on the other hand, since PromptAgent runs iterative optimization upon evaluator GPT-4 and generator GPT-3.5, resulting in model-oriented prompt optimization. We thus observe obvious degradation on out-of-box Llama2-7B and Tulu2-7B models. In contrast, our instruction-oriented optimization remains efficacy, and leading the average performances on two small generators.
Case Analysis. In Table 5, we present an example from the downstream GSM8K benchmark for case analysis, which is same as the example shown in Figure 2. We first notice that either Raw Task Response or Ground Truth Response provide correct answers. Nevertheless, the Raw Task Response improperly denotes “list [12,34,56,75]” as a Python data structure. Similarly, both Optimized Prompt generated from GPT APIs provide high-quality paraphrases of the Raw Task Instruction, yet the positive one is superior with chain-of-thought plans (shown with blue in the upper part). We then compare prompt optimization results from our fine-tuned models. The SFT strategy fails with incoherent, misleading clarifications (shown with red in the middle). In contrast, all preference fine-tuned models provide more coherent and concise generations. Particularly, two IPO-based optimizer manages to provide chain-of-thought plans (shown with blue in the bottom), and IPL-IPO-70B further provides cheating notes (noted with underlines).

### v2

@@ 3 Experiments
The ultimate target of FIPO lies in the general performance enhancement with downstream generators $\bm{M}_{g}$ (Eq. 3). Herein, in the evaluation, we first use fine-tuned optimizer $\bm{M}_{o}$ to produce optimized testing prompts $\bm{\hat{x}}_{t}^{o}$, then obtain optimized testing response $\bm{\hat{y}}_{t}^{o}$ and naive testing response $\bm{\hat{y}}^{n}_{t}$ for answer quality checking, as shown in the bottom right corner of Figure 3. We begin with our experimental settings (§ 3.1), efficacy presentation and comparisons against online ad-hoc APO methods (§ 3.2.1), then follow with analysis of different fine-tuning strategies (§ 3.2.2), and case analysis (§ 3.2.3) [^5] Hyperparameters and training cost are in Appendix B..
@@ 3.1.1 Baselines
We compare FIPO with two SOTA APO methods: APE Zhou et al. (2023) and PromptAgent Wang et al. (2023a). APE, stands for Automatic Prompt Engineer, is a template-based strategy that ask one LLM to generate a pool of candidate prompts based on the templates, then select one according to the evaluation scores. PromptAgent eliminates templates and replaces with Monte Carlo Tree Search Abramson (2014) for using a evaluator model to guide the generator. Both APE and PromptAgent are training-free, aiming to model-oriented APO in an ad-hoc manner, while we realize completely offline training. Following former works, we use GPT-3.5-turbo as the in-box generator and GPT-4 as the optimizer in both baselines.
3.1.2 Foundation models
We use Tulu2 models as our bases, which is a fine-tuned version of Llama2 Touvron et al. (2023) trained on a mix of publicly available datasets Ivison et al. (2023). We fine-tuning local optimizer with Tulu2-13B and Tulu2-70B.
Generation Multi-choice
Llama2-7B Naive Prompt 8.89 31.21 62.78 43.09 46.58 41.73
Touvron et al. (2023) FIPO Optimizer 11.70 33.50 69.37 52.11 54.56 48.10
Tulu2-13B Naive Prompt 39.06 36.49 76.62 55.13 57.43 52.53
Ivison et al. (2023) FIPO Optimizer 40.17 40.26 78.58 57.68 59.10 54.79
Baichuan2-13B Naive Prompt 46.81 37.95 68.56 51.88 57.46 52.36
Yang et al. (2023a) FIPO Optimizer 48.12 39.95 74.77 56.88 58.32 54.35
Table 2: Evaluation results of various downstream generator LLMs, using the best local optimizer from Table 5.
GPT-3.5-turbo Llama2-7B Tulu2-7B
BBH Tasks Naive APE PromptAgent FIPO Naive APE PromptAgent FIPO Naive PromptAgent FIPO GPT-4
Penguins 0.595 0.747 0.797 0.614 0.436 0.402 0.426 0.458 0.336 0.470 0.525 0.542
Geometry 0.227 0.490 0.670 0.738 0.536 0.499 0.544 0.542 0.089 0.350 0.415 0.397
Epistemic 0.452 0.708 0.806 0.865 0.611 0.680 0.775 0.754 0.581 0.691 0.644 0.603
Object Counting 0.612 0.716 0.860 0.660 0.318 0.347 0.312 0.350 0.690 0.577 0.637 0.674
Temporal 0.720 0.856 0.934 0.756 0.362 0.440 0.643 0.683 0.155 0.176 0.210 0.176
Causal Judgment 0.470 0.570 0.670 0.759 0.547 0.583 0.547 0.616 0.832 0.882 0.891 0.888
Average 0.513 0.681 0.790 0.732 0.468 0.492 0.541 0.567 0.447 0.524 0.554 0.547
Table 3: Comparison between FIPO optimizer, previous methods and GPT-4’s prompt optimization.
3.1.3 Evaluation Benchmarks
We include five benchmarks across two most common formats: (1) GSM8k Cobbe et al. (2021), a generative dataset contains 1.3k primary level math questions; (2) BigBenchHard (BBH) Suzgun et al. (2023), which involves 23 challenging reasoning tasks. BBH has 6.4k testing samples, and asks for generative answering; (3) PiQA Bisk et al. (2020), in which 1.8k common physical knowledge questions are proposed, alongside with multiple candidate choices; (4) CosmosQA Huang et al. (2019). There are around 3k commonsense-based multi-choice questions in CosmosQA, equipped with four candidate options; (5) MMLU Hendrycks et al. (2021), which is one of the largest multi-choice benchmarks. MMLU covers 14k questions. For our FIPO experiments, we report results on all five benchmarks. While differently, since both APE and PromptAgent only provide evaluations on 6 tasks of BBH, we report the comparison results aligned with their settings. As for the result metrics, either “generation” or “multi-choice” benchmarks takes few-shot format with strict answering templates (e.g., “The answer is X”). Herein, we are able to report the accuracy score for all benchmarks.
3.2 Experimental Results
3.2.1 Efficacy of FIPO
General improvements of FIPO. We first check if FIPO-optimized prompts have general gains on different downstream generators in Table 2, and conclude that our optimized instruction-oriented prompts are generally beneficial. The optimized prompts help Llama2-7B, Tulu2-13B and Baichuan2-13B models gain 6.37%, 2.26% and 1.99% performance growth on average. The improvements reduced as the generator grows larger. This may be due to the fact that larger LLMs has richer intrinsic world knowledge, therefore rely less on the enhanced instructions.
Inference Per Test
Dataset Training Fee Time
APE \$0 \$0 \$5 2h
PromptAgent \$0 \$0 \$5 2h
GPT-4 \$0 \$0 \$4 1h
FIPO \$300 10h \$0 30s
Table 4: Cost report of different methods.
Comparable optimization capability against online ad-hoc APO, even better. We would like to compare the local FIPO optimizer with previous methods and direct prompt optimization using GPT-4. Table 3 reports experimental results on six BBH tasks. When tested with in-box generator GPT-3.5-turbo, PromptAgent obtain significant 28% improvements over all tasks, compared with naive prompts. Although our FIPO optimizer does not use testing feedback from GPT-3.5-turbo, we still maintain over 22% optimization performance. On the other hand, when apply two ad-hoc APO methods on out-of-box Llama2-7B and Tulu2-7B models, we find their enhancement shrinks more than 10%. In contrast, FIPO remains efficacy over different testing generators, and even beats GPT-4.
In terms of cost, shown in Table 4, building FIPO optimizer will have \$300 one-time collecting cost and 10 hours training cost, which equals to 75 usages of GPT-4. However, once FIPO models are prepared, we save a lot of time as the optimization can be done locally. FIPO leads to less overall costs when optimizing over thousands of prompts.
@@ 3.2.1 Efficacy of FIPO
Naive Prompt 24.77 36.21 73.35 51.17 51.22 47.79
Best 13B Optimizer 18.42 33.55 72.03 49.14 48.20 44.93
SFT-70B 21.43 32.92 74.39 49.97 51.55 46.96
DPO-70B 27.74 35.56 74.17 54.93 52.73 49.07
IPO-70B 25.00 39.21 76.84 56.01 54.29 50.94
IPL-DPO-70B 25.13 35.25 74.95 50.46 52.12 48.10
IPL-IPO-70B 26.67 39.60 77.11 56.71 56.02 52.13
IPO-70B-gen 22.72 41.91 74.53 53.60 52.74 50.23
IPO-70B-partial 23.08 40.03 76.22 54.29 51.99 49.59
Table 5: Evaluation results of FIPO optimizer fine-tuned with different strategies, tested with same Tulu2-7B model. The number attached with the benchmark is the number of in-context examples (e.g., BBH (3) means 3-shot testing on BBH). FIPO only optimizes the last task instruction, leaving in-context examples remained.
3.2.2 Fine-tuning Analysis
We fine-tune Tulu2-13B and -70B models as our local optimizer through different strategies as mentioned above. We use downstream performance of Tulu2-7B for analyzing the effectiveness of different fine-tuning strategies. Our findings are here:
Small optimizer fails. “Small” Tulu2-13B are not up to the difficult prompt optimization task (the 4th line of Table 5). The average testing scores of using optimized prompts even worse than using the naive prompts written by human.
@@ 3.2.2 Fine-tuning Analysis
SFT vs. DPO/IPO vs. IPL. When simply provide a best optimized prompt as the supervision label, indicated by SFT-70B results, the end-to-end prompt optimization is still a hard task. While when contrastive preference supervisions are provided, there are promising improvements obtains, ranging from marginal 0.31% to significant 4.34%. In terms of different preference fine-tuning methods, IPO beats DPO in either solely fine-tuning, or joint integration in our proposed IPL pipeline, which may due to its regularized design in Eq. 10. We analyze more fine-tuning details and the self-rewarding benefits of IPL in Appendix C and  D.
Dataset diversification is necessary. In the bottom of Table 5, we present ablation studies of preprocessing dataset diversification, mentioned in section 2.4. In specific, IPO-70B-gen stands for not diversifying half of the training set into multi-choice format, which is introduced as type 5,6,7 and 8 in Figure 3. As for IPO-70B-partial, we only use type 3,4,7 and 8 in Figure 3 as pairwise diversification templates. The corrupted optimizer contributes to a decrease on GSM8K and multi-choice benchmarks, while improves the BBH.
Naive Testing Prompt I $\bm{x}^{n}_{t_{1}}$ from BBH:
Jane was born on the last day of February in 2000. Today is her 16-year-old birthday. What is the date a month ago in MM/DD/YYYY?
Optimized Testing Prompt I $\bm{x}^{o}_{t_{1}}$ by FIPO
As today marks Jane’s 16th birthday, determine the date from exactly one month prior. Jane was born on the last day of February in the year 2000, a leap year. Ensure your response is in the format MM/DD/YYYY.
Naive Testing Prompt II $\bm{x}^{n}_{t_{2}}$ from MMLU:
Three of the following are classic security properties; which one is not? A. Confidentiality, B. Availability, C. Correctness, D. Integrity.
Optimized Testing Prompt II $\bm{x}^{o}_{t_{2}}$ by FIPO
Which of the following is NOT one of the core security principles? A. Confidentiality, B. Availability, C. Correctness, D. Integrity. Provide the letter corresponding to the incorrect option.
Naive Testing Prompt III $\bm{x}^{n}_{t_{3}}$ from CosmosQA:
A team of engineers constructing signal lights for airplanes that they can use to guide them to runways are attempting to determine the brightness needed for the pilot to be able to detect the tower at 1 mile away. They set the light to a test brightness and establish communication with an inbound pilot. When the pilot is 1 mile away from the tower, he says he cannot see the light. In terms of Signal Detection Theory, what would this be called?
Optimized Testing Prompt III $\bm{x}^{o}_{t_{3}}$ by FIPO
According to Signal Detection Theory, when a pilot cannot detect a signal at a set brightness level, it is called a ’miss’. In contrast, it is called a ’hit’. A team of engineers is testing the brightness of signal lights for airplanes to guide them to runways. They establish communication with an inbound pilot and set the light to a test brightness. When the pilot is one mile away from the tower, he reports that he cannot see the light. What term from Signal Detection Theory describes this situation?
Naive Testing Prompt IV $\bm{x}^{n}_{t_{4}}$ from GSM8K:
Lily has a list of numbers: [12, 34, 56, 75]. Calculate the average value of the list.
Optimized Testing Prompt IV $\bm{x}^{o}_{t_{4}}$ by FIPO
@@
Table 6: Examples from testing benchmarks.
3.2.3 Case Analysis.
In Table 6, we present several examples from the downstream testing benchmarks, discussing the efficacy and shortcomings of FIPO. Particularly, we smear (key optimized content with blue), and overwhelmed (cheating notes with underlines). The 1st optimized prompt of BBH case explicitly mentions that ”2000 is a leap year”, which is a key detail for calculating dates in February. The 2nd optimized prompt of MMLU question capitalizes “NOT” to draw attention to the negative aspect, ensuring the model focuses on identifying the incorrect option. It also explicitly instructs the model to provide the letter of the incorrect option, reducing ambiguity. And the 3rd optimized prompt of CosmosQA case provides definitions for “miss” and “hit” according to Signal Detection Theory, making it easier for the model to understand and apply the correct term. While in the last GSM8K case, FIPO breaks down the calculation into clear, step-by-step instructions, ensuring the model understands the process of finding the average. However, it provides overwhelmed cheating notes of the final answer. We manually confirm that this issue mostly occurs in GSM8K and BBH’s math calculation questions, happening <10%, in Appendix E.
4 Related Work
Preference Optimization for LLMs allows LLMs to align with human minds in a more nuanced way Pan et al. (2023b), compared to SFT. Proximal Policy Optimization (PPO) is one of the well-known preference optimization approach to first train a reward model with pairwise human preference data, then align LLMs with the reward model via reinforcement learning Ouyang et al. (2022); Bai et al. (2022). Despite its efficacy, PPO is often blamed for its training instability and expensive costs. To this end, Direct Preference Optimization (DPO) have been proposed, aiming to align LLMs through implicit modeling, thereby eliminating the flaws associated with the explicit use of reward models Rafailov et al. (2023). A series of following work of DPO are further presented Azar et al. (2023); Zhao et al. (2023b); Ethayarajh et al. (2024).